OpenStudy (anonymous):

Simplify the expression. x times the square root of the quantity 64 x cubed, plus 2 x times the square root of x, plus 8 times the square root of the quantity 2 x If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (anonymous):

\[x \sqrt{64x^3} + 2x \sqrt{x} + 8\sqrt{2x}\]

OpenStudy (anonymous):

@zimmah can you help me?

OpenStudy (anonymous):

sorry i was busy

OpenStudy (anonymous):

wouldn't that be x?

OpenStudy (anonymous):

oh wait, got to simplify it first, of course it's not just x, it threw me off for a moment

OpenStudy (anonymous):

oh hey its alright

OpenStudy (anonymous):

x*(64x^3)^(1/2) + 2x(x)^1/2 + 8(2x)^1/2

OpenStudy (anonymous):

rewrite 8x^(2/2) *x^(3/2) + 2x^(2/2)*x^(1/2) + 8\(\sqrt{2}\)x^(1/2)

OpenStudy (anonymous):

is that hard to read? because i'm a bit lazy to properly format it, but if it's hard to understand i'll just work it out for you properly

OpenStudy (anonymous):

wait so what would the answer be?

OpenStudy (anonymous):

not on that yet, just checking if you follow the steps i take

OpenStudy (anonymous):

and if you can easily read it this way, or if i should format it properly

OpenStudy (anonymous):

ok sorry im just anxious considering ive been working on this damn algebra for like 2 hours now because im really slow at doing math and i can read it

OpenStudy (anonymous):

alright. The reason why i wrote the x as x^(2/2) instead of just x, is because we have to simplify it and write it in standard form. So 8x^(2/2) *x^(3/2) + 2x^(2/2)*x^(1/2) + 8*√2*x^(1/2) makes 8x^(5/2) + 2x^(3/2) + 8*2^(1/2)*x^(1/2) since x^(5/2) is clearly higher than the others, 8 is the leading coefficient. If i was going to fast don't be ashamed to ask clarification

OpenStudy (anonymous):

sorry for the ???? that's just a copy-pasted square root sign.

OpenStudy (anonymous):

ok haha i was about to ask what that was

OpenStudy (anonymous):

some equation as above, but copypasting messes up the format if there's latex commands in it

OpenStudy (anonymous):

same*

OpenStudy (anonymous):

you see what i did there or is there a step i need to clarify?

OpenStudy (anonymous):

one second im writing it down on paper to see if i understand how you got it because it is a tiny confusing in that format

OpenStudy (anonymous):

\[\Large x(64x^3)^{1/2} + 2x(x)^{1/2} + 8(2x)^{1/2}\] \[\Large 8x^{2/2} *x^{3/2} + 2x^{2/2}*x^{1/2} + 8\sqrt{2x}\] \[\Large8x^{5/2} + 2x^{3/2} + 8\sqrt{2x}\]

OpenStudy (anonymous):

same steps, proper format this time

OpenStudy (anonymous):

ok yea thats way better!

OpenStudy (anonymous):

alright i think i understand how to get it

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

it's not particularly hard, just some technique you need to be aware of :)

OpenStudy (anonymous):

yea well its hard when your really bad at math :p haha

OpenStudy (anonymous):

thank you man for your help

OpenStudy (anonymous):

math takes practice more than anything, and often a good teacher helps a lot

OpenStudy (anonymous):

could you help me on another problem? its my last problem left

OpenStudy (anonymous):

i had the luck to have great math teachers most of my life, so i know a lot about it

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

im going to post it and @ at you

OpenStudy (anonymous):

ok